OpenStudy (anonymous):

A rectangle is rawn in the first quadrant where one of its vertices is at the origin and the opposite vertex is a point on a graph: y=72-2x Determine the coordinates for the vertex of the rectangle that lies on the line y=72-2x that would create the rectangle with the greatest area.. To approach this, can I make y=(x)(72-2x) in vertex form then just find the vertex from that?

OpenStudy (anonymous):



OpenStudy (jim766):

in this situation...all of the points are possible lengths and widths y = 72 -2x x y --------- 1 70 the width would be 1 and 70 would be length area = 70 2 68 2 68 area = 136 just need to find the pair of points that create the biggest area

OpenStudy (jim766):

try vertex 18, 36

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

that would work right @Jim766

OpenStudy (anonymous):

I got (13,338)

OpenStudy (jim766):

that is not a point on the line y = 72-2x 338 = 72 - 2(13) it doesnt equal

OpenStudy (anonymous):

awe poo alright

OpenStudy (jim766):

you have a graphing calc?